OpenStudy (anonymous):

I need to convert 540 degrees into radians. please check my answer

OpenStudy (anonymous):

these are the choices A)1/2 pi B) pi c) 3pi D) 2pi

OpenStudy (anonymous):

i converted the 540 into radians and I got this 540 degrees = 9.424777961

OpenStudy (anonymous):

so does this mean that the answer is 3pi ?

ganeshie8 (ganeshie8):

3pi is right !

OpenStudy (anonymous):

3 x (22/7) = 9.424

ganeshie8 (ganeshie8):

u could have used below : 180 degrees = pi radians

OpenStudy (anonymous):

so I did right , thanks again @ganeshie8

ganeshie8 (ganeshie8):

so, 540 degrees = 540 * pi/180 radians = 3pi radians

OpenStudy (anonymous):

got it thanks :)

ganeshie8 (ganeshie8):

good job !